OpenStudy (anonymous):

solve 2sin(2x)-cot(x)cos(x)=1-csc(x) I have: 2(2 sin(x)cos(x)-cos(x)/sin(x)*cos(x)+1/sin(x)-1 not sure if that is correct yet

OpenStudy (lgbasallote):

just continuing from where you are right now.. \(\Large 2sinxcosx - \frac{cos^2x}{sinx} + \frac{1}{sinx} - 1 = 0\) btw...are you sure this is not proving??

OpenStudy (anonymous):

im sure its solving

OpenStudy (anonymous):

i would multiply by sin x

OpenStudy (lgbasallote):

continuing from that...i'll group some terms...then tell me if you see something familiar (by that i mean if you get an idea on how to solve) \(\Large (2sinxcosx - 1) + (\frac {-cos^2x}{sinx} + \frac{1}{sinx}) = 0\) are you seeing something? or need i go on?

OpenStudy (anonymous):

what happened to the 2 in the beginning?

OpenStudy (lgbasallote):

hmm oh yeah..it's supposed to be 4sinxcosx then the rest

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lemme work it out, i'll show you in a min....if thats ok

OpenStudy (lgbasallote):

sure :DDD

OpenStudy (lgbasallote):

interesting problem btw..needs some imagination haha

OpenStudy (anonymous):

my teacher is nutzzzo

OpenStudy (lgbasallote):

lol =))) so...figured it out yet?

OpenStudy (anonymous):

should i multiply by sin x?

OpenStudy (lgbasallote):

multiply what by sinx?

OpenStudy (anonymous):

the whole thing

OpenStudy (anonymous):

get rid of the fractions

OpenStudy (lgbasallote):

hmm...it works on rational fractions....go on..try

OpenStudy (anonymous):

ok, so this is what i think i should have: \[\sin(x)(\sin(x)-1)=0 and \cos(x)+1=0\]

OpenStudy (lgbasallote):

huh? why'd you separate 2 equations?

OpenStudy (anonymous):

x=0 x=π x=π/2 x=2π/3